# Curvature and local matchings of conference graphs and extensions

### v2

@@ 2.2 Strongly/Amply regular graph
If $b=1$, then $v_{1}$ has no neighbor in $N_{y}$, that is, $x_{1}=0$. Thus by Lemma 3.3, $v_{1}$ has $\beta-1$ neighbors in $\Delta_{xy}$ and $d-\alpha-1$ neighbors in $P_{xy}$. If condition (2) or (3) or (4) holds, then $\beta-1>\alpha=|\Delta_{xy}|$, which is a contradiction. If condition (5) holds, then $|P_{xy}|=n-2d+\alpha\leq\frac{1}{2}(d-\alpha-1)<d-\alpha-1$, which is a contradiction. If condition (1) holds, we have $\beta-1=\alpha$ and hence $v_{1}$ is adjacent to all vertices in $\Delta_{xy}$. Since $n<3d-2\alpha$, we have $|P_{xy}|=n-2d+\alpha\leq d-\alpha-1$. Thus, $|P_{xy}|=d-\alpha-1$ and $v_{1}$ is adjacent to all vertices in $P_{xy}$. Therefore, $v_{1}$ has in total $d-1$ neighbors in $\Delta_{xy}\cup P_{xy}$. Hence, $v_{1}$ has no neighbors in $N_{x}\cup N_{y}$.
@@
$2\displaystyle 2q=d-\alpha-$ 2 . (3.5)
Thus, $d-\alpha-1$ is odd, which is a contradiction.
@@
Let $X=x_{1}+\cdots+x_{b}$, then $X$ is the number of edges between $S:=\{v_{1},\ldots,v_{b}\}$ and $T$.
@@
Since $n-2d+\alpha\leq d-\alpha-1$ under each condition $(1)-(5)$, it follows by Cauchy inequality that
@@
Set $N_{x}=\{v_{1},\ldots,v_{d-\alpha-1}\}$, and let $D_{k}$ be the number of neighbors of $v_{k}$ in $S$, where $11\leq k\leq d-\alpha-$ 1. Counting the number of ordered pairs $(v_{k},v_{j})$ with $v_{k}\in N_{x}$, $v_{j}\in S$ and $v_{k}v_{j}\in E$, we have
@@
We divide the remaining argument of Case 2 into four subcases.
@@
(ii) Suppose that the condition (2) holds. Then we derive from (3) that
$M_{1}\geq\frac{1}{2}\left[\frac{\big{(}(\alpha+1)b-X\big{)}^{2}}{\alpha+1}-\big{(}(\alpha+1)b-X\big{)}\right].$ (3.14)
Inserting (3.14), (3.8), (3.9), and (3) into (3), we have
$\left(\frac{1}{\alpha+1}+\frac{2}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}$ $-\left(4b+\frac{2b}{d-\alpha-1}\right)X$
$+\frac{b^{2}}{d-\alpha-1}+b+(b^{2}-b)(d-\alpha-1)\leq 0.$
$b\left(2+\frac{1}{d-\alpha-1}\right)^{2}$ $\geq\left[\frac{1}{\alpha+1}+\frac{2}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b}{d-\alpha-1}+1+(b-1)(d-\alpha-1)\right]$
$>2(b-1)+d-\alpha-1+\frac{b}{(d-\alpha-1)^{2}}+\frac{2}{d-\alpha-1}+\frac{(b-1)(d-\alpha-1)}{\alpha+1}.$
That is,
$0>d-\alpha-2+\frac{2}{d-\alpha-1}-\frac{d}{\alpha+1}-\left(3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}\right)b.$ (3.15)
Since $d\leq 2\beta$, we obtain
$3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}>3-\frac{d}{\alpha+1}\geq 3-\frac{2\beta}{\alpha+1}\geq 0,$
and hence
00 0 $>d-\alpha-2+\frac{2}{d-\alpha-1}-\frac{d}{\alpha+1}-\left(3+\frac{4}{d-\alpha-1}-\frac{d}{\alpha+1}\right)\frac{d-\alpha}{2}$
$=d-\alpha-4-\frac{3(d-\alpha)}{2}+\frac{d(d-\alpha-2)}{2(\alpha+1)}.$ (3.16)
$d<\alpha+\sqrt{10\alpha+\frac{41}{4}}+\frac{3}{2},$
which is a contradiction.
(iii) Suppose that either condition (3) or (4) holds. Inserting (3), (3.8), (3.9) and $M_{4}\geq 0$ into the inequality (3) leads to
@@
$b\left(\frac{\beta-1}{\alpha}+1\right)^{2}>\left[\frac{1}{\alpha}+\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
@@
$\frac{2b(\beta-1)}{\alpha}+2b>$ $1+\frac{bd-b\beta-d+1}{\alpha}$
$+\left[\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
@@
$>$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+\left(\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right)b(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
$>$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)+\frac{b(\beta-1)}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right).$
@@
$<\left[\frac{2(\beta-1)}{\alpha}+1-\frac{d-\beta}{\alpha}-\frac{\beta-1}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)\right]b.$
@@
Therefore, we obtain
@@
$\left|d-\frac{3}{2}\beta\right|<\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+4\alpha+24\beta-20}).$
@@
(iv) Suppose that the condition (5) holds. Then we have $n-2d+\alpha\leq\frac{1}{2}(d-\alpha-1)$. It follows from (3.7) that
$M_{2}\geq\frac{1}{2}\left[\frac{2\left((d-\alpha-1)b-X\right)^{2}}{d-\alpha-1}-\left((d-\alpha-1)b-X\right)\right].$ (3.18)
Inserting (3), (3.18), (3.9) and (3) into the inequality (3), we have
$\left(\frac{1}{\alpha}+\frac{3}{d-\alpha-1}+\frac{1}{b-1}\right)X^{2}-\left(\frac{2b(\beta-1)}{\alpha}+4b-\frac{2(\alpha-\beta+1)b}{d-\alpha-1}\right)X$
$\hskip 19.91684pt+\frac{b^{2}(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b^{2}+\frac{(\alpha-\beta+1)^{2}b^{2}}{d-\alpha-1}-(d-\alpha)b-(\alpha-1)b^{2}\leq 0.$ (3.19)
Since the discriminant of the above quadratic polynomial is non-negative, we deduce
$b\left(\frac{\beta-1}{\alpha}+2-\frac{\alpha-\beta+1}{d-\alpha-1}\right)^{2}\geq\left(\frac{1}{\alpha}+\frac{3}{d-\alpha-1}+\frac{1}{b-1}\right)\times$
$\hskip 19.91684pt\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right).$ (3.20)
It is clear that
$\frac{1}{\alpha}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
$>$ $\frac{b(\beta-1)^{2}}{\alpha^{2}}+\frac{2b(d-\alpha-1)}{\alpha}-\frac{d-\alpha+\alpha b-b}{\alpha}.$ (3.21)
Since $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}$, we have
$d-\alpha-1\geq\alpha+2-\frac{(\beta-1)^{2}}{\alpha}>\alpha-\beta+3\geq 3.$ (3.22)
Thus, we estimate
$\frac{3}{d-\alpha-1}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
$=$ $\frac{3b(\beta-1)^{2}}{(d-\alpha-1)\alpha}+6b+\frac{3(\alpha-\beta+1)^{2}b}{(d-\alpha-1)^{2}}-3-\frac{3}{d-\alpha-1}-\frac{3(\alpha-1)b}{d-\alpha-1}$
$>$ $\frac{3b(\beta-1)^{2}}{(d-\alpha-1)\alpha}+6b+\frac{(\alpha-\beta+1)^{2}b}{(d-\alpha-1)^{2}}-4-\frac{3(\alpha-1)b}{d-\alpha-1}.$ (3.23)
By the condition that $d\geq 2\alpha+3-\frac{(\beta-1)^{2}}{\alpha}$, we derive
$\frac{1}{b-1}\left(\frac{b(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)b+\frac{(\alpha-\beta+1)^{2}b}{d-\alpha-1}-(d-\alpha)-(\alpha-1)b\right)$
$=$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(\alpha-1)$
$\hskip 39.83368pt+\frac{1}{b-1}\left(\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(d-1)\right)$
$>$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{(\alpha-\beta+1)^{2}}{d-\alpha-1}-(\alpha-1)$
$>$ $\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-(\alpha-1),$ (3.24)
where the last inequality above follows from Lemma 3.2. Inserting (3.21),(3),(3) into (3.20) yields
$-\frac{d}{\alpha}+\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-\alpha-2$
$<$ $\left(\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\right)b.$ (3.25)
By Lemma 3.2, we have $\frac{3}{2}\beta\geq d$. By (3.22), we have $d-\alpha-1\geq\alpha-\beta+4$. Thus, the coefficient of $b$ satisfies
$\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\geq\frac{\beta-3}{\alpha}-1+0+1\geq 0.$
It follows from (3) that
$-\frac{d}{\alpha}+\frac{(\beta-1)^{2}}{\alpha}+2(d-\alpha-1)+\frac{2(\alpha-\beta+1)^{2}}{\alpha}-\alpha-2$
$<$ $\left(\frac{4\beta-2d-3}{\alpha}-\frac{\alpha-\beta+4}{d-\alpha-1}+\frac{(\beta-1)(\alpha-\beta+1)}{\alpha(d-\alpha-1)}+1\right)\frac{d-\alpha}{2}$
$\leq$ $\frac{(4\beta-2d-3)(d-\alpha)}{2\alpha}-\frac{\alpha-\beta+4}{2}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha(d-\alpha-1)}+\frac{d-\alpha}{2}$
$<$ $\frac{(4\beta-2d-3)(d-\alpha)}{2\alpha}-\frac{\alpha-\beta+4}{2}+\frac{(\beta-1)(\alpha-\beta+1)}{2\alpha}+\frac{1}{2}+\frac{d-\alpha}{2}.$ (3.26)
That is,
$02d^{2}+(\alpha-4\beta+1)d-6\alpha\beta+\alpha+7\beta^{2}-14\beta+7<$ 0 .
Since $06\beta-d-1>$ 0, we have
$\alpha>\frac{2d^{2}-4\beta d+7\beta^{2}+d-14\beta+7}{6\beta-d-1},$
which is a contradiction.
@@
In addition to the cases in Theorem 1.4, our method can deal with many other cases. Here, we comment on some possible ways to extend our conclusion.
@@
    Suppose that there are $e$ edges within $S:=\{v_{1},\ldots,v_{b}\}$ and $f$ edges between $S$ and $N_{x}\backslash S$. Note that $\sum_{k=1}^{b}D_{k}=2e$ and $\sum_{k=b+1}^{d-\alpha-1}D_{k}=f$. The inequality (3) can be improved as
@@
To illustrate the above remark, we show that the Lin–Lu–Yau curvature of a particular strongly regular graph with $n\geq 3d-2\alpha$ achieves the upper bound $\frac{2+\alpha}{d}$.
@@
Consider a strongly regular graph $G$ with parameters (324, 152, 70, 72) derived from regular symmetric Hadamard matrices with constant diagonal (RSHCD), which is constructed in [11, Section 5.3.2] based on [18, Lemma 11]. Note that the condition $n<3d-2\alpha$ does not hold. Inserting the parameters into (3), (3.7), (3.9), (3), and (3), we have
@@
where $412\leq b\leq\frac{d-\alpha}{2}=$ 41. Considering the discriminant, we have
@@
It is direct to check that the inequality above can not hold for $2\leq b\leq 41$. Therefore, the Lin–Lu–Yau curvature of $G$ is
@@
Now, let’s prove Theorem 1.3. Recall that, in the setting of Theorem 1.3, we have $(n,d,\alpha,\beta)=(4\gamma+1,2\gamma,\gamma-1,\gamma)$. By the condition $(1)$ of Theorem 1.4, we only need to verify the cases that $\gamma$ is equal to $2$, $3$, $4$, $5$, or $6$.
@@
In case that $X=5$, we derive from (3), (3.8) and (3.9) that $M_{1}\geq 3$, $M_{2}\geq 5$ and $M_{3}\geq 4$ respectively. We claim that $M_{4}\geq 1$. Set $N_{x}=\{v_{1},\ldots,v_{5}\}$. Denote by $e$ the number of edges within $\{v_{1},\ldots,v_{5}\}$ and $f$ edges between $\{v_{1},\ldots,v_{5}\}$ and $N_{x}\backslash\{v_{1},\ldots,v_{5}\}$. Observe that $X=2e+f$ by Lemma 3.3 (ii). Thus, the number of edges within $\{v_{1},v_{2},v_{3}\}$ is at most $2$. If there are 2 edges within $\{v_{1},v_{2},v_{3}\}$, then $M_{4}\geq 1$, as claimed. If the number of edges within $\{v_{1},v_{2},v_{3}\}$ is at most $1$, then there are at least $3$ edges between $\{v_{1},v_{2},v_{3}\}$ and $\{v_{4},v_{5}\}$, which implies that either $v_{4}$ or $v_{5}$ is adjacent to two vertices of $\{v_{1},v_{2},v_{3}\}$. That is, we have $M_{4}\geq 1$ as claimed.
@@
Suppose that $\gamma=6$, then $b\leq\frac{7}{2}$. This contradicts to is contradictory to $b>\gamma-3$.
We have completed the proof. ∎
To conclude this section, we prove the number theoretic result Corollary 1.5.
@@
Consider the corresponding Paley graph $P(n)$. By Theorem 1.1 and Theorem 1.3, there is a perfect matching $\mathcal{M}$ between $N_{x}$ and $N_{y}$. Since $|S|\geq\frac{3}{4}(n-1)$, we have $\left|\left(GF(n)\setminus\{x,y\}\right)\setminus S\right|\leq\frac{1}{4}(n-1)-1$. Recall a Paley graph $P(n)$ with $n=4\gamma+1$ is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$. Therefore, we have $|N_{x}|=\gamma=\frac{1}{4}(n-1)$. That is, we have $\left|\left(GF(n)\setminus\{x,y\}\right)\setminus S\right|\leq|N_{x}|-1$. Then it is clear that $GF(n)\backslash S$ is not a vertex cover of $\mathcal{M}$. Suppose that $wz\in\mathcal{M}$ is not covered by $GF(n)\setminus S$, then the result follows. ∎
4 Proof of Theorem 1.7 and Corollary 1.8
In this section, we prove Theorem 1.7 and Corollary 1.8. We first recall the following facts about eigenvalues of strongly regular graphs, see, e.g., [17, Page 220].
Theorem 4.1.
Let $G=(V,E)$ be a strongly regular graph with parameters $(n,d,\alpha,\beta)$. Then the Laplacian eigenvalues of $G$ are
$\lambda_{1}=0,$
$\lambda_{2}=1-\frac{(\alpha-\beta)+\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}{2d},$
$\lambda_{3}=1-\frac{(\alpha-\beta)-\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}{2d}.$
The corresponding multiplicities are
$m_{1}=1,$
$m_{2}=\frac{1}{2}\biggl{(}(n-1)-\frac{2d+(n-1)(\alpha-\beta)}{\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}\biggr{)},$
$m_{3}=\frac{1}{2}\biggl{(}(n-1)+\frac{2d+(n-1)(\alpha-\beta)}{\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}}\biggr{)}.$
By Theorem 1.1, Theorem 1.7 is a direct consequence of the following result.
Theorem 4.2.
Let $G=(V,E)$ be a strongly regular graph with parameters $(n,d,\alpha,\beta)$ such that $\beta\geq\alpha$. If $\lambda_{2}\leq\frac{2+\alpha}{d}$, then the equality holds, i.e., $\lambda_{2}=\frac{2+\alpha}{d}$.
Proof.
Since $\lambda_{2}\leq\frac{2+\alpha}{d}$, we have by Theorem 4.1 that
$2d-3\alpha+\beta-4\leq\sqrt{(\alpha-\beta)^{2}+4(d-\beta)}.$ (4.1)
Since $G$ is connected and not complete, we have $d\geq\alpha+2$. Therefore, we derive
$2d-3\alpha+\beta-4\geq 2(\alpha+2)-3\alpha+\beta-4\geq 0$
In the second inequality above, we use our assumption $\beta\geq\alpha$. It follows that we can square both sides of the inequality (4.1). Hence, we have
$(2d-3\alpha+\beta-4)^{2}\leq(\alpha-\beta)^{2}+4(d-\beta).$
The above inequality can be rearranged as
$4(d-\alpha-1)(d-2\alpha+\beta-4)\leq 0$
Since $d\geq\alpha+2$, we deduce
$d\leq 2\alpha-\beta+4.$ (4.2)
On one hand, since $\beta\geq\alpha$, we have $d\leq\alpha+4$. That is, $\alpha+2\leq d\leq\alpha+4$. On the other hand, since $d\geq\beta$, we have $\beta\leq\alpha+2$.
Observe that the multiplicity numbers $m_{2}$ and $m_{3}$ in Theorem 4.1 must be positive integers. Hence, $(\alpha-\beta)^{2}+4(d-\beta)$ has to be a perfect square. Then, it is straightforward to check that there are only three possible parameters $(n,d,\alpha,\beta)$ for $G$:
  • (a)
    $(n,\alpha+4,\alpha,\alpha)$;
  • (b)
    $(n,\alpha+3,\alpha,\alpha+1)$;
  • (c)
    $(n,\alpha+2,\alpha,\alpha+2)$.
Notice that $n$ can be uniquely determined by the other parameters $d,\alpha,\beta$. Indeed, we have by Theorem 2.6 that
$n-d-1=\frac{d(d-\alpha-1)}{\beta}.$ (4.3)
In the case of $(n,\alpha+4,\alpha,\alpha)$. Since $\frac{d(d-\alpha-1)}{\beta}=3+\frac{12}{\alpha}$ must be a positive integer, the possible values of $\alpha$ are $1$, $2$, $3$, $4$, $6$, $12$. By (4.3), we have $n=\alpha+\frac{12}{\alpha}+8$ and hence the multiplicity numbers $m_{2}=\frac{(4+\alpha)(6+\alpha)}{4\alpha}$ and $m_{3}=\frac{3(2+\alpha)(4+\alpha)}{4\alpha}$. Since $m_{2}$ and $m_{3}$ have to be positive integers, the possible values of $\alpha$ are $2$, $4$, $6$, $12$. Therefore, the possible parameters in this case are
  • (1)
    $(16,6,2,2)$;
  • (2)
    $(15,8,4,4)$;
  • (3)
    $(16,10,6,6)$;
  • (4)
    $(21,16,12,12)$.
In the case of $(n,\alpha+3,\alpha,\alpha+1)$. Since $\frac{d(d-\alpha-1)}{\beta}=2+\frac{4}{\alpha+1}$ must be a positive integer, the possible values of $\alpha$ are 00, $1$, $3$. Therefore, the possible parameters are
  • (5)
    $(10,3,0,1)$;
  • (6)
    $(9,4,1,2)$;
  • (7)
    $(10,6,3,4)$.
In the case of $(n,\alpha+2,\alpha,\alpha+2)$. By (4.3) and Theorem 4.1, we calculate that $m_{2}=\frac{4+\alpha}{2}$ and $m_{3}=\frac{2+\alpha}{2}$. Since $m_{2}$ and $m_{3}$ have to be positice integers, $\alpha$ must be even. Therefore, the possible parameters are
  • (8)
    $(2k,2k-2,2k-4,2k-2),k\geq 2$.
By Theorem 4.1, it is direct to check for all possible graphs $(1)-(8)$ that
$\lambda_{2}=\frac{2+\alpha}{d}.$
This completes the proof. ∎
Proof of Theorem 1.7.
Since $G$ is Lichnerowicz sharp, we have
$\lambda_{2}=\min_{xy\in E}\kappa(x,y)\leq\frac{2+\alpha}{d},$
where the last inequality is due to Theorem 1.1. By Theorem 4.2, we conclude that $\min_{xy\in E}\kappa(x,y)=\lambda_{2}=\frac{2+\alpha}{d}$. Therefore, we obtain $\kappa(x,y)=\frac{2+\alpha}{d}$ for all $xy\in E$ by applying Theorem 1.1 again. This completes the proof. ∎
We recall the following classificaiton result from [14].
Theorem 4.3.
[14, Theorems 6.3 and 6.4] The Lichnerowicz sharp strongly regular graphs with smallest adjacency eigenvalue $-2$ are precisely the following ones: The cocktail party graphs $CP(n)$, $n\geq 2$, the lattice graphs $L_{2}(n)\cong K_{n}\times K_{n}$, $n\geq 3$, the triangular graphs $T(n)\cong J(n,2)$, $n\geq 5$, the demi-cube $Q^{5}_{(2)}$, and the Schläfli graph.
Proof of Corollary 1.8.
By Theorem 4.1, any strongly regular graph with $\lambda_{2}=\frac{2+\alpha}{d}$ has the smallest adjacency eigenvalue $-2$. Thus, by Theorem 4.3, we determine all the Lichnerowicz sharp strongly regular graphs with $\beta\geq\alpha$ as following: The cocktail party graphs $CP(k),k\geq 2$, the Paley graph $P(9)$, the Johnson graphs $J(5,2)$ and $J(6,2)$ and the demi-cube $Q^{5}_{(2)}$. ∎

### v3

@@ 2.2 Strongly/Amply regular graph
If $b=1$, then $v_{1}$ has no neighbors in $N_{y}$, that is, $x_{1}=0$. Thus, by Lemma 3.3, $v_{1}$ has $\beta-1$ neighbors in $\Delta_{xy}$ and $d-\alpha-1$ neighbors in $P_{xy}$. If condition (2) or (3) holds, then $\beta-1>\alpha=|\Delta_{xy}|$, which is a contradiction. If condition (1) holds, we have $\beta-1=\alpha$ and hence $v_{1}$ is adjacent to all vertices in $\Delta_{xy}$. Since $n<3d-2\alpha$, we have $|P_{xy}|=n-2d+\alpha\leq d-\alpha-1$. Thus, $|P_{xy}|=d-\alpha-1$ and $v_{1}$ is adjacent to all vertices in $P_{xy}$. Therefore, $v_{1}$ has in total d1d-1 neighbors in $\Delta_{xy}\cup P_{xy}$. Hence, $v_{1}$ has no neighbors in $N_{x}\cup N_{y}$.
@@
$2q=d-\alpha-2.$ (3.5)
Thus, $d-\alpha-1$ is odd, contradicting (3.4).
@@
Let $X=x_{1}+\cdots+x_{b}$, then $X$ is the number of edges between $S$ and $T$.
@@
Since $n-2d+\alpha\leq d-\alpha-1$ under each condition (1)(5)(1)-(5), it follows by Cauchy inequality that
@@
Set $N_{x}=\{v_{1},\ldots,v_{d-\alpha-1}\}$, and let $D_{k}$ be the number of neighbors of $v_{k}$ in $S$, where $1\leq k\leq d-\alpha-1$. Counting the number of ordered pairs $(v_{k},v_{j})$ with $v_{k}\in N_{x}$, $v_{j}\in S$ and $v_{k}v_{j}\in E$, we have
@@
We divide the remaining argument of Case 2 into two subcases.
@@
(ii) Suppose that either condition (2) or (3) holds. Inserting (3), (3.8), (3.9) and $M_{4}\geq 0$ into the inequality (3) leads to
@@
$b\left(\frac{\beta-1}{\alpha}+1\right)^{2}\geq\left[\frac{1}{\alpha}+\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
@@
$\frac{2b(\beta-1)}{\alpha}+2b$
$\geq$ $1+\frac{bd-b\beta-d+1}{\alpha}+\left[\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right]\left[\frac{b(\beta-1)^{2}}{\alpha}-(\beta-1)+(b-1)(d-\alpha-\beta)\right].$
@@
$\geq$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+\left(\frac{1}{d-\alpha-1}+\frac{1}{b-1}\right)b(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right)$
$>$ $\frac{bd-b\beta-d+1}{\alpha}+d-\alpha-1+\frac{b(\beta-1)}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)+(\beta-1)\left(\frac{\beta-1}{\alpha}-1\right).$
@@
$<$ $\left[\frac{2(\beta-1)}{\alpha}+1-\frac{d-\beta}{\alpha}-\frac{\beta-1}{d-\alpha-1}\left(\frac{\beta-1}{\alpha}-1\right)\right]b.$ (3.14)
@@
Therefore, by the assumption that $b\leq\frac{d-\alpha}{2}$, the inequality (3.14) yields
@@
$\left|d-\frac{3}{2}\beta\right|<\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+4\alpha+24\beta-20})<\frac{1}{2}\sqrt{4\alpha^{2}-3\beta^{2}+28\beta}).$
@@
In addition to the cases in Theorem 1.3 and [9, Theorem 1.4], our method can deal with many other cases. Here, we comment on some possible ways to extend our conclusion.
@@
[…]
@@
To illustrate the above remark (i), we show that the Lin–Lu–Yau curvature of a particular strongly regular graph with $n\geq 3d-2\alpha$ achieves the upper bound $\frac{2+\alpha}{d}$.
@@
Consider a strongly regular graph $G$ with parameters (324, 152, 70, 72) derived from regular symmetric Hadamard matrices with constant diagonal (RSHCD), which is constructed in [11, Section 5.3.2] based on [16, Lemma 11]. Note that the condition $n<3d-2\alpha$ does not hold. Inserting the parameters into (3), (3.7), (3.9), (3), and (3), we have
@@
where $2\leq b\leq\frac{d-\alpha}{2}=41$. Considering the discriminant, we have
@@
It is direct to check that the inequality above does not hold for $2\leq b\leq 41$. Therefore, the Lin–Lu–Yau curvature of $G$ is
@@
Now, let us prove Theorem 1.4. Recall that, in the setting of Theorem 1.4, we have $(n,d,\alpha,\beta)=(4\gamma+1,2\gamma,\gamma-1,\gamma)$. By the condition $(1)$ of Theorem 1.3, we only need to verify the cases that $\gamma$ is equal to $2$, 33, 44, 55, or $6$.
@@
[…]
@@
Suppose that $\gamma=6$, then $b\leq\frac{7}{2}$, which is contradictory to $b>\gamma-3$. ∎
To conclude this section, we prove the number-theoretic Corollary 1.5.
@@
Consider the corresponding Paley graph $P(n)$. By Theorem 1.1 and Theorem 1.4, there is a perfect matching $\mathcal{M}$ between $N_{x}$ and $N_{y}$. Since $|S|\geq\frac{3}{4}(n-1)$, we have $\left|\left(GF(n)\setminus\{x,y\}\right)\setminus S\right|\leq\frac{1}{4}(n-1)-1$. Recall that a Paley graph $P(n)$ with $n=4\gamma+1$ is a strongly regular graph with parameters $(4\gamma+1,2\gamma,\gamma-1,\gamma)$. Therefore, we have $|N_{x}|=\gamma=\frac{1}{4}(n-1)$. That is, we have $\left|\left(GF(n)\setminus\{x,y\}\right)\setminus S\right|\leq|N_{x}|-1$. Then, it is clear that there exists an edge $wz\in\mathcal{M}$, which is not adjacent to any vertex in $GF(n)\setminus S$, and the desired result follows. ∎